# A unified model for long-period radio transients and white dwarf binary pulsars

### v1

@@ Abstract
In any case, the emission mechanism must differ substantially from the common radio pulsar curvature radiation process, a testable prediction of which is the beam opening angle confined by the last open field lines at the pole. Assuming that the emission site is within the orbital radius ($r_{em}\leq A\leq 2.55R_{\odot}$) places a generous limit on the half-opening angle of $\rho\leq 15^{\circ}$, inconsistent with the best-fit half-opening angle $W_{\text{spin}}=65\pm 3^{\circ}$. Further separating GPM J1839--10 from neutron stars is the lack of PA sweep typically associated with the rotating vector model, although this is not a unifying feature of WD pulsars [26].
@@
Future work will include applying the model to the growing number of LPTs, which will not only test it, but serve to constrain the distribution of their geometries. Of particular interest is the 6.45 h LPT ASKAP J183950.5--075635.0 [27] which has timing variation and pulse profile evolution phenomenologically similar to GPM J1839--10 but with an interpulse, extending the geometric parameter space. Also of interest would be J1755--2527, recently confirmed to be a 1.16 hour intermittent LPT [28], the emission-state-switching 54 minute LPT ASKAP J1935+2148 [29] and CHIME J0630$+$25 [30], whose 421 s period makes it an interesting counterpart to J1912--44.
@@
Our results demonstrate that LPTs like GPM J1839--10 can be understood within the same emission geometry framework as radio emitting WD binary pulsars such as J1912--44 and AR Sco. The diversity in the observed periods, pulse morphologies, and polarisation states arises from variations in orbital synchronisation and viewing geometry. This implies that coherent radio emission from magnetic WD binaries is more common and varied than previously thought, potentially forming a new class of compact binary radio emitters. However, the possible evolutionary connection between these systems remains debated, and we cannot exclude them being at different evolutionary stages but with common properties and geometries allowing them to emit bright polarized radio emission. With enough samples, we expect their evolutionary stages to emerge, and the tools developed in this work will enable insight into binary WD and magnetic field evolution in a way difficult to select from optical surveys alone.
@@ Acknowledgements
Cs.H performed the orbital timing, spin/beat period analysis, polarisation analysis, developed the geometric model, and contributed to the discussion in the manuscript. N.R contributed to the formulation of the model and its physical interpretation, and contributed to the discussion in the manuscript. N.H.-W led the proposals for the radio data toward GPM J1839--10, imaged and produced dynamic spectra from the radio data used in this work, and contributed to the formulation of the model and discussion in the manuscript. S.M contributed to the formulation of the model, and contributed to the discussion in the manuscript. R.P generated the observing scripts and calibrated the data for the VLA observations. E.L processed the ASKAP data and contributed to the polarisation analysis.

### v2

@@ 3 Discussion
In any case, the emission mechanism must differ substantially from the common radio pulsar curvature radiation process, a testable prediction of which is the beam opening angle confined by the last open field lines at the pole. Assuming that the emission site is within the orbital radius ($r_{em}\leq A\leq 2.55R_{\odot}$) places a generous limit on the half-opening angle of $\rho\leq 15^{\circ}$, inconsistent with the best-fit half-opening angle $W_{\text{spin}}=65\pm 3^{\circ}$.
3.1 Polarisation
The emission of GPM J1839-10 is consistently highly linearly polarised (up to 100%) with occasional 90 jumps in polarisation angle (PA) to orthogonal polarisation modes (OPMs). However, Figure 1 reveals that the two pulse groups which appear at different orbital and spin phases are dominated by different OPMs, with a similar amount of variation in each group ($\sigma_{\text{PA}}\approx{\sim}10^{\circ}$). The PAs of individual pulses evolve more gradually and systematically across spin phase than can be seen in the point clusters in Figure 1 (see [19]), but we focus on the average behaviour. Both the single pulse PA variations and the occasional OPM jumps within a pulse group may be due to turbulence in the intrabinary wind, R1.10]but why one OPM should be preferred at a particular spin/orbit phase is unclear. Other spectral characteristics (dispersion measure (DM), rotation measure (RM), spectral shape) remain constant across the orbit, although DM and RM variations can be suppressed by the surrounding magnetospheric plasma.
It can further be seen that there is a small downturn in the PAs at around spin phase 0.05-0.05 (the trailing edge of the first group), as well as a similar upturn in the PAs at spin phase $0.05$ (the leading edge of the second group). Despite the possibility that dim pulses arriving between the two groups could potentially connect the PAs smoothly across intermediate spin–orbit phases (with greater sensitivity), the flatness of the PAs of the two groups themselves, as well as the occasional $90^{\circ}$ jump within each group, lead us to prefer an OPM interpretation over a rotating vector model[34].
Although not well understood, OPMs in pulsars are thought to be magnetospheric in origin, possibly due to birefringent properties of the co-rotating plasma[35]. We advance a similarly broad interpretation here: R1.10]the linear polarisation is dominated by the average magnetic properties of the plasma through which the WD beam propagates along the LOS, switching the visible polarisation mode with spin phase. If the mechanism behind the pulse group PA separation and the short duration OPMs are related, then further study of GPM J1839-10 may give insights into OPMs more broadly.
4 Conclusion
@@
Future work will include applying the model to the growing number of LPTs, which will not only test it, but serve to constrain the distribution of their geometries. Of particular interest is the P=P=6.45-h LPT ASKAP J183950.5-075635.0 [36] which has timing variation and pulse profile evolution phenomenologically similar to GPM J1839-10 but with an interpulse, extending the geometric parameter space. Also of interest would be ASKAP J1755-2527, recently confirmed to be a 1.16-hour intermittent LPT [37], the emission-state-switching 54 minute LPT ASKAP J1935+2148 [38] and CHIME J0630++25 [39], whose 421 s period makes it an interesting counterpart to J1912-44.
Our results demonstrate that LPTs like GPM J1839-10 can be understood within the same emission geometry framework as radio emitting WD binary pulsars such as J1912-44 and AR Sco. The diversity in the observed periods, pulse morphologies, and polarisation states arise from variations in orbital synchronisation and viewing geometry.
@@
The possible evolutionary connection between these systems remains debated, and we cannot exclude them being at different evolutionary stages but with common properties and geometries allowing them to emit bright polarised radio emission. With enough samples, we expect their evolutionary stages to emerge, and the methods developed in this work will enable insight into binary WD and magnetic field evolution in a way difficult to select from optical surveys alone.
@@ Acknowledgements
Cs.H performed the orbital timing, spin/beat period analysis, polarisation analysis, developed the geometric model, and contributed to the discussion in the manuscript. N.R contributed to the formulation of the model and its physical interpretation, and contributed to the discussion in the manuscript. N.H.-W led the proposals for the radio data toward GPM J1839-10, imaged and produced dynamic spectra from the radio data used in this work, and contributed to the formulation of the model and discussion in the manuscript. S.M contributed to the formulation of the model, and contributed to the discussion in the manuscript. R.P generated the observing scripts and calibrated the data for the VLA observations. E.L processed the ASKAP data and contributed to the polarisation analysis.